2 kg of ice at -20degree centigrade is mixed with 5 kg of water at 20degree centigrade in an insulating vessel having a negligible heat capacity. The final mass of water in the vessel. ( the specific heat of water and ice are 1k cal/kg degree centigrade and 0.5 k cal/kg degree centigrade respectively ant the latent heat of fusion of ice is 80 k cal/kg) is
 
1) 7kg  2) 6kg 3) 4kg 4) 2kg

Grade:12th pass

1 Answers

Arun
5 years ago
5 kg of water = 5000 g of water 
Heat lost by water when its temperature drops to 0°C from 20°C 
= 5000 * 20 = 100,000 cals. 
Heat gained by 2000 g of ice in reaching 0°C from - 20°C 
= 2000 * 0.5 * 20 = 20,000 cals. 
Remaining heat from water = 100,000 - 20,000 = 80,000 will melt m gram of ice 
=> m * 80 = 80,000  
=> m = 1000 g. 
=> final mass of water in the container 
= 5000 + 1000 g 
= 6 kg.